### v4

@@ Abstract
Machine learning is playing an increasing role in hydrology, supplementing or replacing physics-based models. One notable example is the use of recurrent neural networks (RNNs) for forecasting streamflow given observed precipitation and geographic characteristics. Training of such a model over the continental United States (CONUS) demonstrated that a single set of model parameters can be used across independent catchments, and that RNNs can outperform physics-based models. In this work, we take a next step and study the performance of RNNs for river routing in land surface models (LSMs). Instead of observed precipitation, the LSM-RNN uses instantaneous runoff calculated from physics-based models as an input. We train the model with data from river basins spanning the globe and test using historical streamflow measurements. The model demonstrates skill at generalization across basins (predicting streamflow in catchments not used in training) and across time (predicting streamflow during years not used in training). We compare the predictions from the LSM-RNN to an existing physics-based model calibrated with a similar dataset and find that the LSM-RNN outperforms the physics-based model. Our results show that RNNs are effective for global streamflow prediction from runoff inputs and motivate the development of complete routing models that can capture nested sub-basis connections.
@@
The surface water cycle is a key component of the climate system (Oki and Kanae, 2006), and river routing of runoff from the land to the ocean is an important transport process simulated in Earth System Models (ESMs) (Li et al., 2015). Routing models provide a freshwater source to ocean models and have a range of additional applications, covering water resource management (He et al., 2017) to flood hazard assessments under climate change scenarios (Wobus et al., 2017).
River basins, the areas drained by individual rivers and their tributaries, tile the land surface into weakly connected domains, with water transport between basins given by the river streamflow. Two main processes are typically considered in river modeling: hillslope routing and river channel routing (Mizukami et al., 2016). Hillslope routing describes how water moves across the landscape, considering factors such as topography, soil characteristics, and vegetation. It is the process that addresses the time lag between instantaneously generated runoff on land and the aggregation of runoff at a river channel, contributing to streamflow. This process is typically unresolved in the river models used in LSMs and must be parameterized. In contrast, river channel routing describes how water moves from upstream channels to downstream outlets within the river network itself. Both the hillslope and river channel routing processes occur simultaneously within a basin.
@@
A physics-based approach to river modeling can be implemented considering different processes at different scales. Lohmann et al. (1996) was one of the first river routing schemes for LSMs, using a linear model to account for hillslope routing in each grid and the linearized Saint-Venant equation to model water transport in-between grids. Later models, like the storage-based schemes of Oki and Sud (1998) and Branstetter (2001), used simplified 1D (one dimensional) kinematic wave routing (KWR) equations to route water for both hillslope and river channel routing indistinguishably, but explicitly accounting for spatial distribution of streamflow under various approximations. More advanced approaches (e.g., Ye et al., 2012; Wu et al., 2014), apply 1D KWR explicitly accounting for the differences between hillslope and river channel routing. Moreover, Li et al. (2013) additionally accounted for tributary channels explicitly. More recent models, such as Mizukami et al. (2016), offer flexibility by coupling different possibilities for hillslope and river channel routing in a domain that can be either vector- or grid-based. Physics-based representation of these processes present several advantages. First, the physical equations describing the flow naturally conserve water mass, which is crucial for systems simulated for long periods of time, as is the case in LSMs. Second, the interpretability of the model is straightforward since one knows exactly what physical laws are being used. Finally, physical laws have a long history of success in physical modeling; thus, these approaches are commonly employed in streamflow forecasting models across disciplines. Despite these advantages, however, physical models tend to perform poorly at regional and global basin generalizability, and it has been argued that this is due to challenges in expressing routing processes across scales and locations using simple physical laws (Nearing et al., 2021). For a more detailed overview of physics-based routing models, see Shaad (2018).
@@
Recently, a class of recurrent deep learning models, referred to as Long-Short-Term-Memory (LSTM) models, have outperformed physics-based models on the rainfall-runoff problem (Kratzert et al., 2019b). In Kratzert et al. (2018), an LSTM was trained to model the entire land hydrology system for the CONUS. The model took observed precipitation as input (along with other dynamic inputs such as near-surface temperature, surface pressure, etc.) and then simulated streamflow at the outlet of gauged catchments, implicitly modeling snowpack, soil storage, runoff, hillslope routing, and river channel routing. In Kratzert et al. (2019b), attributes such as topography, vegetation, and soil properties from different catchments were added to the model to improve its performance. The model did not explicitly account for routing between basins and treated each catchment independently. Nonetheless, the model showed good performance in regional calibrations, where a single set of parameters is learned using data from multiple catchments at once. Further studies also showed that these types of models were successful at basin generalization, predicting streamflow on the outlet of catchments not included in the training set (Kratzert et al., 2019a) and more recently, similar models have demonstrated greater global basin generalizability than physics-based models (Nearing et al., 2024). In particular, these results suggest that models can represent the unresolved processes in hillslope and channel routing with accurately. A drawback to using these machine learning (ML) models in LSMs is that they are not guaranteed to conserve mass a priori (or produce otherwise physical output, such as positive streamflow); however, constraints can be enforced by adapting the architecture of the network, for example, to be mass conserving (Hoedt et al., 2021). More importantly, though machine-learning based models have been used for routing between basins in specific regions (Moshe et al., 2020), these types of models have not been used yet for this purpose over entire continents, which would also be a necessary step in order to implement them in LSMs.
@@ 0.1 Our contributions
Our goal in this work is to explore the use of LSTMs for river routing in a global LSM. To do so, we make multiple alterations to the rainfall-runoff model (an LSTM) of Kratzert et al. (2019b), including in model architecture and training and validation procedures, as follows.
@@
    We construct a globally consistent dataset to train and validate runoff-driven models. We incorporate runoff variables from reanalysis and use a globally unified system of basin characterizations, to ensure that our routing model can be integrated into an LSM in the future. This also requires addressing consistency between gauged catchments and geographical definitions of the nested sub-basins provided by our base dataset (Lehner et al., 2008). A similar idea was first introduced by the Caravan dataset (Kratzert et al., 2023), using precipitation instead of runoff and catchments instead of globally consistent sub-basins.
@@
    We consider the mass conservation properties of the LSTM (in Appendix D).
This work hence represents a first step toward using LSTMs for river routing in LSMs, but we still treat each basin as independent as in Kratzert et al. (2019b). The problem of routing between basins is left for future work.
@@ 0.2 Outline and Notation
Regarding notation, we wish to be explicit about the following. We note that the terms catchment and basin are often used interchangeably in the hydrology literature; however, in this paper we use “basin” (and sub-basin) for the units of topographical subdivision of the world into drainage areas and “catchment” for the drainage area of specific gauges. Additionally, we use the term generalization in time as described above (generalizing to time periods unseen in calibration, but only for basins used in calibration) in place of what the hydrology community often refers to as performance in gauged basins. We use the term basin generalization as described above (generalizing to basins not used in calibration) in place of what the hydrology community refers to as regionalization or performance in ungauged basins. This is because whether a basin is gauged or not is a property of the basin, while time and basin generalizability are experimental design concepts. For example, only gauged basins can be used for validating a model, regardless of if that model is calibrated in a “gauged" (generalization in time) or “ungauged" (basin generalization) fashion).
The paper is structured as follows. Section 1 (Methods) presents the various components of our model, including data engineering and the training of the ML model. Section 2 (Results) presents our findings regarding time and basin generalization, a comparison to a physics-based model, and an investigation of model performance by basin attributes. Section 3 (Discussion and conclusion) summarizes the main results and outlines future directions, including applying the LSTM to routing between basins.
@@ 1.1.2 Dynamic inputs
Figure 1: (a) Illustration of the Monte-Carlo algorithm used for transforming gridded ERA5-Land data into basin-specific data for HydroSHEDS basin 2070017000 (shaded pink). Grid cells that are completely inside the basin have a weight of 1 because their entire area lies within the basin. Grid cells outside the basin have weights less than 1, representing the fraction of their area within the basin. Other basins are shaded grey, and their boundaries are outlined in black. The white area is the sea (in this case, Mar des Baleares on the east coast of Spain). (b) GRDC gauges and the river network structure within the same basin. In the figure, the chosen gauge (white circle) is the one used in the calibration of our model for this specific basin, as it better represents the entire drainage area of the basin. The other gauge has a smaller catchment area, so it represents a smaller fraction of the behavior of the basin.
@@ 1.1.3 Streamflow
The resulting runoff and streamflow timeseries on the basin in Figures 1a and 1b is shown in Figure 2.
Figure 2: Surface, sub-surface runoff and streamflow timeseries for basin 2070017000 of HydroSHEDS.
@@ 2.2.1 Comparison with a physics-based model
Below, we compare the results of GloFAS to the LSTM in both a time- and basin-split experiment. However, several caveats are necessary to make. First, we do not know the exact time-split used in the GloFAS calibration. This means that the validation time frame of the LISFLOOD model may coincide with the test period used in our comparison here, potentially simplifying the prediction task of the time-split experiment for GloFAS relative to the LSTM. Second, the locations of only GRDC gauges that were used to calibrate GloFAS are known and other datasets may have been used in place of the same gauges (GloFAS team members, private communication). This allows for a basin-split experiment comparison, but (a) some of the gauges in our test set may have been used in the calibration of GloFAS, and (b) the unknown time-split may mean that for some of the test basins, GloFAS is generalizing in both space and time. The former makes the GloFAS prediction task easier than the LSTM task, while the latter makes the GloFAS prediction task harder than the LSTM task, since the LSTM is trained and tested in different basins but using the same time window. For a large enough temporal training period and for a global set of basins, we would expect that generalizing in space and time is similar to generalizing in space, since that is the much more challenging task. For all of these reasons, we note the basin- and time-split experiments with a "*" in the Table and Figure presented in this Section.
@@ 2.2.2 Comparison with other LSTM models
In this section, we present simulated timeseries along with observations for various values of NSE and KGE to get a visual sense of performance. For four different basins, Figure 6 shows the predictions of the LSTM model in the global time-split configuration, the GloFAS reanalysis, and the observed streamflow at the corresponding GRDC gauge.
In the first example (Figure 6a), we have a poor performance in both NSE/KGE scores for both the LSTM model ($0.03$/$-0.16$) and for GloFAS reanalysis ($0.14$/$-0.15$). For both models, this is due to underprediction of a peak in the streamflow. This specific basin (1061638580 in HydroSHEDS level 06) is located in South Africa, and the corresponding gauge is located at the Seekoei River. It has a significantly lower streamflow than the other basins in the figure. (We come back to trends in the model performance with basin characteristics in the following section.) The next example (Figure 6b) lies around the median performance of our model: ($0.61$/$0.63$) for the LSTM model and ($0.61$/$0.62$) for GloFAS reanalysis. This basin (6050344660 in HydroSHEDS level 05) is located in Brazil, and the corresponding gauge is located at the Itacaiúnas River. The third plot (Figure 6c) shows an example where the LSTM model has a very good performance ($0.71$/$0.82$) and outperforms GloFAS slightly ($0.62$/$0.80$). This specific basin (7070250410 in HydroSHEDS level 07) is located in Canada, and the corresponding gauge is located at the North French River. The last plot (Figure 6d) shows a case where GloFAS ($0.84$/$0.91$) outperforms the LSTM model ($0.78$/$0.88$). This specific basin (7060363050 in HydroSHEDS level 06) is located at the border between Canada and the United States, and the corresponding gauge is located at the Saint John River.
Figure 6: Timeseries simulated by the LSTM model (pink), by GloFAS reanalysis (green), and the observed timeseries from GRDC gauges (blue) in four different basins. The gauges chosen are described in more detail in Section 2.3. The legend in (b) equally applies in (a), (c) and (d).
@@ 2.4 Geographic patterns
We also investigate the aridity index, which is provided in the HydroATLAS dataset (Linke et al., 2019) following the dataset of Zomer et al. (2008). The aridity index is given by the ratio between the mean annual precipitation and the mean annual evapotranspiration, and is calculated on a per grid cell basis. Hence, lower values of aridity index represent arid climates and higher values represent humid climates, for further discussion, one may reference to the newest version of the dataset in Zomer et al. (2022). In our analysis, the aridity index has been found to be a strong predictor of the model’s performance. Figure 8 shows a trend toward poorer performance in drier regions (i.e., regions with lower aridity index), consistent with findings in Feng et al. (2020), but now on a global scale. (For context, the basin depicted in Figure 6a has an aridity index of 0.26.) This trend extends to other variables, such as mean runoff, for which the model also shows poorer performance in drier regions.
@@ 3 Discussion and conclusion
In that effort, we have trained and validated an LSTM at the task of predicting streamflow from modeled runoff worldwide. We began our analysis by contrasting the results from an reanalysis precipitation-driven model, which holistically represents land hydrology, with those of a modeled runoff-driven model. For the United States, we found that an LSTM trained only to route runoff performs comparably to one designed to simulate the entire land hydrology system, even when trained on potentially less accurate and biased runoff data. Our exploration into the model’s generalization capabilities revealed that a globally trained model, as opposed to one trained exclusively using data from the USA, achieved superior performance when both were fed runoff data in a time-split configuration. This improvement underscores the potential benefits of incorporating diverse global data. However, when evaluating models trained in the basin-split configuration, the performance gain was not as pronounced, highlighting the complex challenge of global basin generalization. This challenge is exacerbated by imbalances in global data availability and the wide range of possible climate conditions across the globe, which may not be well sampled in geographically localized training data (e.g., from the USA).
@@
Additionally, our analysis revealed a correlation between the model’s performance and the aridity index. While streamflow in arid basins can be modeled well by the LSTM, it is also true that all basins with a poor NSE have a lower aridity index. This suggests that drier regions pose challenges for the LSTM model, but that other basin features may affect performance as well. The dependence on aridity could be due to the model’s difficulty in capturing short-term precipitation events that trigger streamflow peaks in these areas, as seen in Figure 6a (Feng et al., 2020). Moreover, the ERA5-Land reanalysis is driven with precipitation from ERA5, known to have biases in the tropics (Lavers et al., 2022), which could lead to biases in the runoff of ERA5-Land in these regions. As a consequence, our river model could be learning to correct these biases in addition to routing water. This is a possible pitfall for machine learning models trained with model output and not with observed data. The LSTM model also underperformed in regions likely underrepresented in the training data. Despite these challenges, the model’s time generalizability was consistent across different basin sizes, even if better basin generalizability was observed for bigger basins, as discussed in Appendix C. We highlight that the model does not need to be re-calibrated for the evaluation of different basin levels, which correspond to different spatial resolutions of LSMs.
@@
We also took steps to compare our model performance with the performance of others. In general, this is challenging due to differences in experimental design, loss function choice, and data used. These differences are motivated by differences in intended application of the model, and hence these comparisons are not always meaningful to make. That said, we would like to highlight two main conclusions we drew from our comparisons.
@@
Using a similar dataset, experimental design, and LSTM architecture, we found that a model trained with reanalysis data (our USA time-split precipitation model) had a significantly worse performance compared to the LSTM of Kratzert et al. (2019b) which was trained with observations (a drop in median NSE from 0.73 to 0.52). This indicates that modeled variables (e.g., precipitation or runoff from reanalysis) suffers from larger errors than observed precipitation, which is not surprising, and that these errors affect the quality of the predicted streamflow. An LSTM trained using modeled runoff - for which observations are not possible - may learn both to predict streamflow and to correct biases in modeled runoff. However, for our use case of a river model routing water within an LSM, and not modeling the entire land hydrology system, this may be unavoidable.
@@
In contrast, the physics-based LISFLOOD model (GloFAS) is more similar to our model in that it uses modeled runoff (calculated internally) as a dynamic input to perform streamflow routing. We have compared the models on a subset of the gauges used to calibrate GloFAS (time-split) and on a subset of gauges not used for calibration (basin-split). The LSTM model showed more accurate simulations in both scenarios, but note that there are still some differences in experimental design between the two. Unfortunately, a recalibration of the model using our data and experimental design is beyond the scope of this work – the reported calibration time of GloFAS is of the order of weeks (GloFAS team members, private communication).
@@
In order to integrate our model with an LSM, it must be extended to channel routing and it must respect physical priniciples like mass conservation. To our knowledge, no article has demonstrated a machine learning-based approach for global routing simulations, although some approaches show how to account for inter-basin connections in a restricted region (Moshe et al., 2020). On the other hand, physical models such as LISFLOOD model routing between basins and hence are currently the standard for LSMs. One may interpret this work as the hillslope routing component of a complete routing model. In that case, two possible ways to extend this to route water between basins are to use a physics-based channel routing model or to use similar neural networks to additionally perform river channel routing. For the latter, designing training strategies to mitigate error accumulation across connected basins may be necessary. Moreover, incorporating architectural adjustments to ensure mass conservation (e.g., Hoedt et al., 2021), is a necessary step for coupling with climate models, potentially enhancing long-term simulation accuracy, although this might not be necessary for integration into LSMs used by weather forecasting models. In conclusion, we have demonstrated that machine learning based river models exhibit basin and time generalizability, a requirement for use in a global climate model. Our study presents the first step toward using these models in the LSM component of climate models as well as for short, medium and long range hydrological forecasting using runoff from LSMs within weather forecasting models. The results presented here motivate further research to extend these models for comprehensive river routing.

### v5

@@ Abstract
Machine learning is playing an increasing role in hydrology, supplementing or replacing physics-based models. One notable example is the use of recurrent neural networks (RNNs) for forecasting streamflow given observed precipitation and geographic characteristics. Training of such a model over the continental United States (CONUS) has demonstrated that a single set of model parameters can be used across independent catchments, and that RNNs can outperform physics-based models. In this work, we take a next step and study the performance of RNNs for river routing in land surface models (LSMs). Instead of observed precipitation, the LSM-RNN uses instantaneous runoff calculated from physics-based models as an input. We train the model with data from river basins spanning the globe and test it using historical streamflow measurements. The model demonstrates skill at generalization across basins (predicting streamflow in catchments not used in training) and across time (predicting streamflow during years not used in training). We compare the predictions from the LSM-RNN to an existing physics-based model calibrated with a similar dataset and find that the LSM-RNN outperforms the physics-based model: a gain in median NSE from 0.56 to 0.64 (time-split experiment) and from 0.30 to 0.34 (basin-split experiment). Our results show that RNNs are effective for global streamflow prediction from runoff inputs and motivate the development of complete routing models that can capture nested sub-basis connections.
@@
The surface water cycle is a key component of the climate system (Oki and Kanae, 2006), and river routing of runoff from the land to the ocean is an important transport process simulated in land surface models (LSMs) within climate models (Li et al., 2015). Routing models provide a freshwater source to ocean models and have a range of additional applications, covering water resource management (He et al., 2017) to flood hazard assessments under climate change scenarios (Wobus et al., 2017).
River basins, the areas drained by individual rivers and their tributaries, tile the land surface into weakly connected domains, with water transport between basins given by the river streamflow. Two main processes are typically considered in river modeling: hillslope routing and river channel routing (Mizukami et al., 2016). Hillslope routing describes how water moves across the landscape, considering factors such as topography, soil characteristics, and vegetation. It is the process that leads to the time lag between instantaneously generated runoff on land and the aggregation of runoff at a river channel, contributing to streamflow. This process is unresolved in the river models used in LSMs and must be parameterized. In contrast, river channel routing describes how water moves from upstream channels to downstream outlets within the river network itself. Both the hillslope and river channel routing processes occur simultaneously within a basin.
@@
A physics-based approach to river modeling can be implemented considering different processes at different scales. Lohmann et al. (1996) was one of the first river routing schemes for LSMs, using a linear model to account for hillslope routing in each grid and the linearized Saint-Venant equation to model water transport in between grids. Later models, like the storage-based schemes of Oki and Sud (1998) and Branstetter (2001), used simplified 1D kinematic wave routing (KWR) equations to route water for both hillslope and river channel routing jointly, but explicitly accounting for the spatial distribution of streamflow under various approximations. More advanced approaches (e.g., Ye et al., 2012; Wu et al., 2014), apply 1D KWR explicitly accounting for the differences between hillslope and river channel routing. Moreover, Li et al. (2013) additionally accounted for tributary channels explicitly. More recent models, such as Mizukami et al. (2016), offer flexibility by coupling different possibilities for hillslope and river channel routing in a domain that can be either vector- or grid-based. Physics-based representation of these processes present several advantages. First, the physical equations describing the flow naturally conserve water mass, which is crucial for systems simulated for long periods of time, as is the case in LSMs. Second, the interpretability of the model is straightforward since one knows exactly what physical laws are being used. Finally, physical laws have a long history of success in physical modeling; thus, these approaches are commonly employed in streamflow forecasting models across disciplines. Despite these advantages, however, physical models tend to perform poorly at regional and global basin generalizability, and it has been argued that this is due to challenges in expressing routing processes across scales and locations using simple physical laws (Nearing et al., 2021). For a more detailed overview of physics-based routing models, see Shaad (2018).
@@
Recently, a class of recurrent deep learning models, referred to as Long-Short-Term-Memory (LSTM) models, have outperformed physics-based models on the rainfall-runoff problem (Kratzert et al., 2019b). In Kratzert et al. (2018), an LSTM was trained to model the entire land hydrology system for the CONUS. The model took observed precipitation as input (along with other dynamic inputs such as near-surface temperature, surface pressure, etc.) and then simulated streamflow at the outlet of gauged catchments, implicitly modeling snowpack, soil storage, runoff, hillslope routing, and river channel routing. In Kratzert et al. (2019b), attributes such as topography, vegetation, and soil properties from different catchments were added to the model to improve its performance. The model did not explicitly account for routing between basins and treated each catchment independently. Nonetheless, the model showed good performance in regional calibrations, where a single set of parameters is learned using data from multiple catchments at once. Further studies also showed that these types of models were successful at basin generalization, predicting streamflow on the outlet of catchments not included in the training set (Kratzert et al., 2019a); more recently, similar models have demonstrated greater global basin generalizability than physics-based models (Nearing et al., 2024). In particular, these results suggest that models can represent the unresolved processes in hillslope and channel routing accurately. A drawback to using these machine learning (ML) models in LSMs is that they are not guaranteed to conserve mass a priori (or produce otherwise physical output, such as positive streamflow); however, constraints can be enforced by adapting the architecture of the network, for example, to be mass conserving (Hoedt et al., 2021). More importantly, though machine-learning based models have been used for routing between basins in specific regions (Moshe et al., 2020), these types of models have not been used yet for this purpose over entire continents, which is a necessary step in order to implement them in LSMs.
@@ 0.1 Our contributions
Our goal in this work is to explore the use of LSTMs for river routing in a global LSM. To do so, we make multiple alterations to the rainfall-runoff LSTM of Kratzert et al. (2019b), including in model architecture and training and validation procedures:
@@
    We construct a globally consistent dataset to train and validate runoff-driven models. We incorporate runoff variables from reanalysis and use a globally unified system of basin characterizations, to ensure that our routing model can be integrated into an LSM in the future. This also requires consistency between gauged catchments and geographical definitions of the nested sub-basins provided by our base dataset (Lehner et al., 2008). A similar idea was first introduced by the Caravan dataset (Kratzert et al., 2023), using precipitation instead of runoff and catchments instead of globally consistent sub-basins.
@@
    We consider the mass conservation properties of the LSTM (Appendix D).
This work represents a first step toward using LSTMs for river routing in LSMs; however, we still treat each basin as independent, as in Kratzert et al. (2019b). The problem of routing between basins is left for future work.
@@ 0.2 Outline and Notation
The terms catchment and basin are often used interchangeably in the hydrology literature. In this paper, we use “basin” (and sub-basin) for the units of topographical subdivision of the world into drainage areas, and “catchment” for the drainage area of specific gauges. Additionally, we use the term generalization in time as described above (generalizing to time periods unseen in calibration, but only for basins used in calibration), in place of what the hydrology community often refers to as performance in gauged basins. We use the term basin generalization as described above (generalizing to basins not used in calibration), in place of what the hydrology community refers to as regionalization or performance in ungauged basins. We choose this terminology because whether a basin is gauged or not is a property of the basin, while time and basin generalizability are experimental design concepts. For example, only gauged basins can be used for validating a model, regardless of whether that model is calibrated in a “gauged" (generalization in time) or “ungauged" (basin generalization) fashion.
The paper is structured as follows. Section 1 presents the various components of our model, including data engineering and the training of the ML model. Section 2 presents our findings regarding time and basin generalization, a comparison to a physics-based model, and an investigation of model performance by basin attributes. Section 3 summarizes the main results and outlines future directions, including applying the LSTM to routing between basins.
@@ 1.1.2 Dynamic inputs
Figure 1: (a) Illustration of the Monte-Carlo algorithm used for transforming gridded ERA5-Land data into basin-specific data for HydroSHEDS basin 2070017000 (shaded pink, located at the east coast of Spain). Grid cells that are completely inside the basin have a weight of 1 because their entire area lies within the basin. Grid cells outside the basin have weights less than 1, representing the fraction of their area within the basin. Other basins are shaded grey, and their boundaries are outlined in black. The white area is the sea (in this case, Mar des Baleares). (b) GRDC gauges and the river network structure within the same basin. In the figure, the chosen gauge (white circle) is the one used in the calibration of our model for this specific basin, as it better represents the entire drainage area of the basin. The other gauge has a smaller catchment area, so it represents a smaller fraction of the behavior of the basin.
@@ 1.1.3 Streamflow
The resulting runoff and streamflow timeseries on the basin in Figure 1 is shown in Figure 2.
Figure 2: Surface, sub-surface runoff, and streamflow during one year for basin 2070017000 of HydroSHEDS, located at the east coast of Spain.
@@ 2.2.1 Comparison with a physics-based model
Below, we compare the results of GloFAS to the LSTM in both a time- and basin-split experiment. However, several caveats should be noted. First, a sensible time-split experiment is challenging, as GloFAS draws different dates for training, validating and testing for each basin. This means that the validation time frame of the LISFLOOD model may coincide with the test period used in our comparison here, potentially simplifying the prediction task of the time-split experiment for GloFAS relative to the LSTM. Second, the locations of only GRDC gauges that were used to calibrate GloFAS are known and other datasets may have been used in place of the same gauges (GloFAS team members, private communication). This allows for a basin-split experiment comparison, but (a) some of the gauges in our test set may have been used in the calibration of GloFAS, and (b) the unknown time-split may mean that for some of the test basins, GloFAS is generalizing in both space and time. The former makes the GloFAS prediction task easier than the LSTM task, while the latter makes the GloFAS prediction task harder than the LSTM task, since the LSTM is trained and tested in different basins but using the same time window. For a large enough temporal training period and for a global set of basins, we would expect that generalizing in space and time is similar to generalizing in space, since that is the much more challenging task. For all of these reasons, we note the basin- and time-split experiments with a "*" in the Table and Figure presented in this Section.
@@ 2.2.2 Comparison with other LSTM models
In this section, we present simulated timeseries along with observations for various values of NSE and KGE to get a visual sense of performance. For four different basins, Figure 6 shows the predictions of the LSTM model in the global time-split configuration, the GloFAS reanalysis, and the observed discharge at the corresponding GRDC gauge.
[…]
Figure 6: Timeseries simulated by the LSTM model (pink), by GloFAS reanalysis (green), and the observed timeseries from GRDC gauges (blue) in four different basins. (a) Basin 6050344660 in HydroSHEDS level 05, located in Brazil, with gauge located at the Itacaiúnas River. (b) Basin 1061638580 in HydroSHEDS level 06, located in South Africa, with gauge located at the Seekoei River. (c) Basin 7050013170 in HydroSHEDS level 05, located in California, with gauge located at the Salinas River. (d) Basin 7060363050 in HydroSHEDS level 06, located at the border between Canada and the United States, with gauge located at the Saint John River.
@@ 2.4 Geographic patterns
We also investigate the aridity index, which is provided in the HydroATLAS dataset (Linke et al., 2019) following the dataset of Zomer et al. (2008). The aridity index is given by the ratio between the mean annual precipitation and the mean annual evapotranspiration, and is calculated on a per grid cell basis. Hence, lower values of aridity index represent arid climates and higher values represent humid climates, for further discussion, one may reference to the newest version of the dataset in Zomer et al. (2022). In our analysis, the aridity index has been found to be a strong predictor of the model’s performance. Figure 8 shows a trend toward poorer performance in drier regions (i.e., regions with lower aridity index), consistent with findings in Feng et al. (2020), but now on a global scale. (For context, the basin depicted in Figure 6b has an aridity index of 0.26.) This trend extends to other variables, such as mean runoff, for which the model also shows poorer performance in drier regions.
@@ 3 Discussion and conclusion
We have trained and validated an LSTM at the task of predicting streamflow from modeled runoff worldwide. We began our analysis by contrasting the results from a precipitation-driven model, which holistically represents land hydrology, with those of a runoff-driven model. For the United States, we found that an LSTM trained only to route runoff performs comparably to one designed to simulate the entire land hydrology system, even when trained on potentially less accurate and biased runoff data. Our investigation of the model’s generalization capabilities revealed that a globally trained model, as opposed to one trained exclusively using data from the USA, achieved superior performance when both were fed runoff data in a time-split configuration. This improvement underscores the potential benefits of incorporating diverse global data. However, when evaluating models trained in the basin-split configuration, the performance gain was not as pronounced, highlighting the complex challenge of global basin generalization. This challenge is exacerbated by imbalances in global data availability and the wide range of possible climate conditions across the globe, which may not be well sampled in geographically localized training data (e.g., from the USA).
@@
Additionally, our analysis revealed a correlation between the model’s performance and the aridity index. While streamflow in arid basins can be modeled well by the LSTM, it is also true that all basins with a poor NSE have a lower aridity index. This suggests that drier regions pose challenges for the LSTM model, but that other basin features may affect performance as well. The dependence on aridity may be due to the model’s difficulty in capturing short-term precipitation events that trigger streamflow peaks in these areas(Feng et al., 2020). Moreover, the ERA5-Land reanalysis is driven with precipitation from ERA5, known to have biases in the tropics (Lavers et al., 2022), which could lead to biases in the runoff of ERA5-Land in these regions. As a consequence, our river model may be learning to correct these biases, in addition to routing water. This is a possible pitfall for machine learning models trained with model output and not with observed data. The LSTM model also underperformed in regions likely underrepresented in the training data, such as in the African continent. Despite these challenges, the model’s time generalizability was consistent across different basin sizes (HydroSHEDS levels), even if better basin generalizability was observed for bigger basins (Appendix C). We highlight that the model does not need to be re-calibrated for the evaluation of different basin levels, which correspond to different spatial resolutions of LSMs.
@@
We also took steps to compare our model performance with the performance of other models. In general, this is challenging due to differences in experimental design, loss function choice, and data used. We highlight two main conclusions we drew from our comparisons:
  1. 1.
@@
    Using a similar dataset, experimental design, and LSTM architecture, we found that a model trained with reanalysis data (our USA time-split precipitation model) had a significantly worse performance (a drop in median NSE from 0.73 to 0.52) compared to the LSTM of Kratzert et al. (2019b), which was trained with observations. This indicates, unsurprisingly, that modeled variables (e.g., precipitation or runoff from reanalysis) suffer from larger errors than observed precipitation, and that these errors affect the quality of the predicted streamflow. An LSTM trained using modeled runoff, for which observations are not available, may learn both to predict streamflow and to correct biases in modeled runoff. Such biases may differ significantly in different regions of the world, and an LSTM provided with terrain particularities and climate characteristics of each basin from static attributes should be capable to use this information to correct runoff model error. This is undesirable, since it prohibits us from using streamflow data to calibrate our runoff models. However, for our use case of a river model routing water within an LSM, and not modeling the entire land hydrology system, this may be unavoidable.
  2. 2.
@@
    The physics-based LISFLOOD model (GloFAS) is more similar to our model in that it uses modeled runoff (calculated internally) as a dynamic input to perform streamflow routing. We have compared the models on a subset of the gauges used to calibrate GloFAS (time-split) and on a subset of gauges not used for calibration (basin-split). The LSTM model showed more accurate simulations in both scenarios: a gain in median NSE from 0.56 to 0.64 (time-split experiment) and from 0.30 to 0.34 (basin-split experiment). Despite these averages differences, GloFAS shows superior fidelity to observations than the LSTM in some cases, while the LSTM is superior in others. However, there are still some differences in experimental design between the two. A recalibration of GloFAS using our data and experimental design for a direct one-on-one comparison is beyond the scope of this work.
@@
In order to integrate our model with an LSM, it must be extended to enable inter-basin channel routing, and it must respect physical principles such as mass conservation. Even if approaches showing how to account for inter-basin connections in a restricted region with ML models have appeared (Moshe et al., 2020), physical models such as LISFLOOD for routing between basins are currently the standard for LSMs. One may interpret our work as the hillslope routing component of a complete routing model. In that case, two possible ways to extend this to route water between basins are to use a physics-based channel routing model, or to use similar neural networks to additionally perform river channel routing. For the latter, designing training strategies to mitigate error accumulation across connected basins may be necessary. Moreover, incorporating architectural adjustments to ensure mass conservation (e.g., Hoedt et al., 2021), is a necessary step for incorporation in climate models.
In conclusion, we have demonstrated that a ML-based river model exhibits basin and time generalizability, a requirement for use in a global climate model. Our study presents the first step toward using such a model in the LSM component of climate models as well as for short-, medium-, and long-range hydrological forecasting using runoff from LSMs within weather forecasting models. The results presented here motivate further research to extend these models for comprehensive river routing.